OpenStudy (anonymous):

The motion of a human body through space can be modeled as the motion of a particle at the body's center of mass as we will study in a later chapter. The components of the displacement of an athlete's center of mass from the beginning to the end of a certain jump are described by the equations xf = 0 + (10.5 m/s)(cos 18.5°)t 0.400 m = 0.610 m + (10.5 m/s)(sin 18.5°)t − ½(9.80 m/s2)t2 where t is in seconds and is the time at which the athlete ends the jump. (a) Identify the athlete's position at the takeoff point. (Let the x and y-direction be along the horizontal and vertical direction resp

OpenStudy (anonymous):

is that xf or f(x)?

OpenStudy (anonymous):

probably f(x)

OpenStudy (anonymous):

you're gonna want to solve for t(0)

OpenStudy (anonymous):

ah ok so its quadratic eq.

OpenStudy (anonymous):

its xf

OpenStudy (anonymous):

its exactly how its written

OpenStudy (anonymous):

from 0.400 m = 0.610 m + (10.5 m/s)(sin 18.5°)t − ½(9.80 m/s2)t^2 we can compute t t1=-0.0581 sec and t2=0.738sec we sub this to xf = 0 + (10.5 m/s)(cos 18.5°)t

OpenStudy (anonymous):

xf = 0 + (10.5 m/s)(cos 18.5°)t xf = 0 + (10.5 m/s)(cos 18.5°)(.738)=7.35meter

OpenStudy (anonymous):

did you get the same answer solving the quadratic eq, for t1 and t2?

OpenStudy (anonymous):

i dont even know how to do it lol thats why i cam on here lol

OpenStudy (anonymous):

ok from 0.400 m = 0.610 m + (10.5 m/s)(sin 18.5°)t − ½(9.80 m/s2)t^2 we can compute t 0=0.610-0.400 + (10.5 m/s)(sin 18.5°)t − ½(9.80 m/s2)t^2 0=0.210 + 3.332t - 4.9 t^2 or 0=0.210 + 3.332t - 4.9 t^2 or -4.9t^2+3.332t+.0210=0 this is a quadratic equation, use the formula for quadratic equation to solve for t where a=-4.9 b=3.332 and c=0.210 sub this to the formula x=[-b-+sqrt(b^2 -4ac)]/2a, x=[-(3.332)-+sqrt((3.332)^2 -4(-4.9)(0.210))]/(2)(-4.9), x1=t1=-0.0581 sec and x2=t2=0.738sec

OpenStudy (anonymous):

ok i read the screenshot, from the formula x=x0+(v0)t-0.5(9.80)t^2 0.400 m = 0.610 m + (10.5 m/s)(sin 18.5°)t − ½(9.80 m/s2)t2 e see that final take off xf=0.400m, initial take off x0=0.610m, initial velocity v0=10.5sin18.5